OpenStudy (anonymous):

3x - 3y = -9 9x - 9y = -27

OpenStudy (imstuck):

What, exactly, are you doing with this?

OpenStudy (anonymous):

3X-3Y+81X-9Y=-27 84X-12Y=-27 28X-4Y=-9 X=\[\left(\begin{matrix}4Y-9 \\ 28\end{matrix}\right)\]

OpenStudy (anonymous):

\[\left[\begin{matrix}?\left(\begin{matrix}4Y-9 \\ 28\end{matrix}\right) & ? \\ ? & ?\end{matrix}\right]\]

OpenStudy (anonymous):

X=(4Y-9)/28